# Advancing Universal Deep Learning for Electronic-Structure Hamiltonian Prediction of Materials

### v1

@@ 1 Introduction
Understanding the electronic structure is fundamental to unraveling how electrons govern the properties of condensed matter systems. This knowledge is essential for predicting a wide range of material characteristics, such as electrical conductivity, magnetism, optical behavior, and chemical activity, which are vital for technologies spanning from electronics to sustainable energy and advanced catalysis. At the heart of these calculations is the challenge of determining the system’s Hamiltonian matrix, whose eigenvalues and eigenstates yield important quantities like energy levels, band structures, and electronic wavefunctions. Traditionally, Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has been the standard approach for these problems. However, as shown in Fig. 1 (a), DFT relies heavily on the self-consistent (SC) procedure, which demands repeated (denoted as $T$ turns), computationally intensive diagonalizations of large matrices—each scaling as $\mathcal{O}(N^{3})$ with system size $N$, making simulations of large or complex materials extremely resource-consuming. Recently, deep learning has emerged as a powerful tool in the physical sciences (Zhang et al., 2025). As shown in Fig. 1 (b), modern deep neural network methods (Gong et al., 2023; Yu et al., 2023; Zhang et al., 2024; Wang et al., 2024b; Li et al., 2025; Yin et al., 2025) can predict Hamiltonians, i.e., the core physical quantities in electronic structure calculations, directly from atomic configurations in an efficient way, circumventing the computationally expensive SC loop and dramatically accelerating computations. This paradigm shift lowers the computational barriers associated with electronic structure calculations, unlocking the simulation and design of unprecedentedly large-scale materials systems, driving new innovation in materials discovery and engineering. Please refer to Appendix A.1 for background introduction.
@@ 2 Related Work
The foundation of deep learning-based electronic-structure Hamiltonian prediction lies in constructing neural networks that respect E(3)-symmetry (with SO(3) as its rotational sub-group), which can be achieved by incorporating group theory-based tensor operators (Geiger & Smidt, 2022). Since traditional non-linear activation functions, when applied directly to SO(3)-equivariant features, may break equivariance, a central research topic is to reconcile strong non-linear expressiveness with strict SO(3)-equivariance. An early attempt to address this problem was the use of gated activation functions (Weiler et al., 2018), which first apply non-linear activations to SO(3)-invariant features and then use them as coefficients to scale the SO(3)-equivariant features. Representative works adopting this mechanism include Equiformer (Liao & Smidt, 2023), eSEN (Fu et al., 2025) for force prediction, and DeepH-E3 (Gong et al., 2023), QHNet (Yu et al., 2023) for Hamiltonian prediction. To further enhance non-linear expressiveness of E(3)-networks, Zitnick et al. (2022) and Passaro & Zitnick (2023) proposed eSCN (equivariant Spherical Channel Networks), which applies non-linear operations to the coefficients obtained from the spherical decomposition of features. This approach has been widely used in tasks such as force-field prediction (Liao et al., 2024) and Hamiltonian prediction (Wang et al., 2024a; b). Nevertheless, eSCN methods project features onto discrete basis functions via inner-product operations, degrading strict SO(3)-equivariance to equivariance with respect to only a discrete sub-group. More recently, Yin et al. (2025) proposed the TraceGrad method, which introduces an SO(3)-invariant supervision signal, i.e., the trace quantity, constructed from Hamiltonian labels, to supervise high-fidelity non-linear SO(3)-invariant features. Then, non-linearity is delivered to SO(3)-equivariant features via a gradient-based mechanism from the SO(3)-invariant ones. This method effectively unifies strict SO(3)-equivariance with strong non-linear expressiveness for Hamiltonian prediction; however, the backbone network it adopted is a simple graph neural network and has not yet evolved into a non-linear equivariant Transformer framework.
Despite these progresses, deep learning methods for Hamiltonian prediction still face substantial challenges on generalization performance, which can be summarized as follows. First, crystalline materials commonly found in nature can be composed of over $65$ different elements from the first six rows of the periodic table, leading to an exceptionally large and heterogeneous input space for deep neural network models. Existing deep learning methods for Hamiltonian prediction typically employ learnable embedding techniques similar to word vectors (Mikolov et al., 2013) to represent nodes (atoms) and edges (atom pairs). These embeddings are randomly initialized and learned directly from the dataset, without incorporating any explicit physical priors. As a result, they struggle to capture the fundamental physical relationships between different atoms and across different material systems, which are crucial for generalization. Second, as illustrated in Figure 1, the regression target, namely the self-consistent electronic-structure Hamiltonian, is inherently high-dimensional and complex, especially when considering SOC effects. For instance, a system containing several tens of atoms may involve nearly several thousands of non-zero Hamiltonian matrix elements that need to be accurately predicted. Most of the existing methods attempt to directly predict the entire self-consistent Hamiltonian matrix, namely $\mathbf{H}^{(T)}$ as formulated in Appendix A.1, placing a heavy burden on the model due to the vast size of the output space, often resulting in optimization difficulties during training and limited generalization to unseen systems. In addition, most existing methods treat the real-space Hamiltonian as the sole regression target, which can lead to sub-optimal physical fidelity in down-stream applications, particularly in capturing low-energy band structures accurately. Although Li et al. (2025) designed a method for molecular systems to reduce the regression space of Hamiltonians and introduced a basis transformation of the Hamiltonian matrix on the loss function to improve the prediction accuracy of down-stream physical quantities, their approach is limited to molecular systems and is not applicable to periodic crystalline materials.
As a result, constructing a unified model that generalizes across diverse crystal prototypes remains challenging, and many existing approaches explicitly constrain their scope. For example, Li et al. (2022), Gong et al. (2023), and Xia et al. (2025) each train and evaluate their methods within a single material system (e.g., MoS2, Bi2Se3, or a-HfO2), without assessing cross-material generalization. More recently, DeepH-2 (Wang et al., 2024b) broadened coverage to systems involving elements primarily from the first four rows of the periodic table; however, they neglected SOC and reduced the orbital basis by omitting $f$-orbitals. While such choices help reduce computational and modeling complexity, they may limit broad applicability to the full diversity of real-world materials. Zhong et al. (2025) developed a Hamiltonian prediction model aimed at a broader range of element types, while also highlighting the challenge of achieving consistently high accuracy across diverse crystal systems. Moreover, open-source datasets with a broad and diverse collection of materials dedicated to training and validating universal Hamiltonian models across the periodic table remain scarce.
To solve these challenges, this work presents an advanced unified deep learning framework together with a large benchmark dataset for Hamiltonian prediction, targeting broader generalization across richer classes of materials.
@@ 3 Method
To handle wider chemical and structural variability while maintaining accuracy and efficiency, we develop an unified Hamiltonian prediction framework along three fronts: (i) input descriptors and output targets, (ii) network architectures, and (iii) training objectives.
@@ A.1 Electronic Structure Calculations: from Density Functional Theory to Deep Learning Methods
To address this challenge, recent approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) have adopted the deep graph learning paradigm to predict the self-consistent Hamiltonians. These methods bypass the iterative and computationally intensive matrix diagonalization steps in traditional DFT algorithms by directly predicting the final converged Hamiltonian matrix $H^{(T)}_{\alpha\beta}$ in a single forward pass. As shown in Eq. (11), the Hamiltonian matrix is inherently sparse: only pairs of atoms within a cutoff radius contribute non-zero elements. Therefore, the total number of Hamiltonian matrix elements that need to be computed scales with the number of local atomic pairs in the system, leading to a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ denotes the average number of neighboring atoms within the cutoff radius per atom. Since the atomic orbital basis functions have finite spatial support, matrix elements vanish beyond a certain inter-atomic distance. In small systems where all atoms lie within each other’s cutoff radius, $\overline{E}\sim N$, and the total number of non-zero elements scales as $\mathcal{O}(N^{2})$. However, in sufficiently large systems, $\overline{E}$ saturates to a constant determined by local geometry, making the number of non-zero Hamiltonian elements scale linearly as $\mathcal{O}(N)$. Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known. Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$ for large systems.
Recently, deep-learning approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) exploit the sparsity of the Hamiltonian, yielding a computational cost that scales approximately linearly with the number of non-zero matrix elements and enabling efficient, scalable prediction of quantum properties in large atomic systems. As a result, they offer a significant efficiency advantage over traditional DFT methods with computational complexity of $\mathcal{O}(TN^{3})$. This efficiency makes them particularly promising for predicting electronic structures of complex atomic systems under limited computational resources, potentially accelerating down-stream application areas like materials simulation and design.

### v2

@@ 1 Introduction
Understanding the electronic structure is fundamental to unraveling how electrons govern the properties of condensed matter systems. This knowledge is essential for predicting a wide range of material characteristics, such as electrical conductivity, magnetism, optical behavior, and chemical activity, which are vital for technologies spanning from electronics to sustainable energy and advanced catalysis. At the heart of these calculations is the challenge of determining the system’s Hamiltonian matrix, whose eigenvalues and eigenstates yield important quantities like energy levels, band structures, and electronic wavefunctions. Traditionally, Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has been the standard approach for these problems. However, as shown in Fig. 1 (a), DFT relies heavily on the self-consistent (SC) procedure, which demands repeated (denoted as $T$ turns), computationally intensive diagonalizations of large matrices—each scaling as $\mathcal{O}(N^{3})$ with system size $N$, making simulations of large or complex materials extremely resource-consuming. Recently, deep learning has emerged as a powerful tool in the physical sciences (Zhang et al., 2025). As shown in Fig. 1 (b), modern deep neural network methods (Gong et al., 2023; Yu et al., 2023b; Zhang et al., 2024; Wang et al., 2024b; Li et al., 2025; Yin et al., 2025) can predict Hamiltonians, i.e., the core physical quantities in electronic structure calculations, directly from atomic configurations in an efficient way, circumventing the computationally expensive SC loop and dramatically accelerating computations. This paradigm shift lowers the computational barriers associated with electronic structure calculations, unlocking the simulation and design of unprecedentedly large-scale materials systems, driving new innovation in materials discovery and engineering. Please refer to Appendix A.1 for background introduction.
@@ 2 Related Work
The foundation of deep learning-based electronic-structure Hamiltonian prediction lies in constructing neural networks that respect E(3)-symmetry, including invariance to translations, as well as invariance and equivariance to rotations (elements of SO(3)) and inversions (Geiger & Smidt, 2022). Since traditional non-linear activation functions, when applied directly to SO(3)-equivariant features, may break equivariance, a central research topic is to reconcile strong non-linear expressiveness with strict SO(3)-equivariance. An early attempt to address this problem was the use of gated activation functions (Weiler et al., 2018), which first apply non-linear activations to SO(3)-invariant features and then use them as coefficients to scale the SO(3)-equivariant features. Representative works adopting this mechanism include DeepH-E3 (Gong et al., 2023) and QHNet (Yu et al., 2023b) for Hamiltonian prediction. To further enhance non-linear expressiveness of E(3)-networks, Zitnick et al. (2022) and Passaro & Zitnick (2023) proposed eSCN (equivariant Spherical Channel Networks), which applies non-linear operations to the coefficients obtained from the spherical decomposition of features. This approach has been widely used in Hamiltonian prediction (Wang et al., 2024a; b; Yu et al., 2025) tasks. Nevertheless, eSCN methods project features onto discrete basis functions via inner-product operations, degrading strict SO(3)-equivariance to equivariance with respect to only a discrete sub-group. Moreover, they replace SO(3) convolutions with SO(2) convolutions, sacrificing strict inversion equivariance. Such compromises may limit the physical soundness of the results. Yin et al. (2025) proposed the TraceGrad method, which introduces an SO(3)-invariant supervision signal, i.e., the trace quantity, constructed from Hamiltonian labels, to supervise high-fidelity non-linear SO(3)-invariant features, and propagates non-linearity to SO(3)-equivariant features via a gradient-based mechanism derived from the SO(3)-invariant ones. This method effectively unifies strict E(3)-symmetry, with strong non-linear expressiveness for Hamiltonian prediction; however, the backbone network it adopted is a simple graph neural network and has not yet evolved into a non-linear equivariant Transformer framework.
@@
Despite these progresses, deep learning methods for Hamiltonian prediction still face substantial challenges on generalization performance, which can be summarized as follows. First, crystalline materials commonly found in nature can be composed of over $65$ different elements from the first six rows of the periodic table, leading to an exceptionally large and heterogeneous input space for deep neural network models. Existing deep learning methods for Hamiltonian prediction typically employ learnable embeddings to represent nodes (atoms) and edges (atom pairs). These embeddings are randomly initialized and learned directly from the dataset, without incorporating any explicit physical priors. As a result, they struggle to capture the fundamental physical relationships between different atoms and across different material systems, which are crucial for generalization. Second, as illustrated in Figure 1, the regression target, namely the self-consistent electronic-structure Hamiltonian, is inherently high-dimensional and complex, especially when considering SOC effects. For instance, a system containing several tens of atoms may involve nearly several thousands of non-zero Hamiltonian matrix elements that need to be accurately predicted. Most of the existing methods attempt to directly predict the entire self-consistent Hamiltonian matrix, namely $\mathbf{H}^{(T)}$ as formulated in Appendix A.1, placing a heavy burden on the model due to the vast size of the output space, often resulting in optimization difficulties during training and limited generalization to unseen systems. In addition, most existing methods treat the real-space Hamiltonian as the sole regression target, which can lead to sub-optimal physical fidelity in down-stream applications, particularly in capturing low-energy band structures accurately. Although Li et al. (2025) designed a method for molecular systems to reduce the regression space of Hamiltonians and introduced a basis transformation of the Hamiltonian matrix on the loss function to improve the prediction accuracy of down-stream physical quantities, their approach is limited to molecular systems and is not applicable to periodic crystalline materials.
As a result, constructing a unified model that generalizes across diverse crystal prototypes remains challenging, and many existing approaches explicitly constrain their scope. For example, Li et al. (2022), Gong et al. (2023), and Xia et al. (2025) each train and evaluate their methods within a single material system (e.g., MoS2, Bi2Se3, or a-HfO2), without assessing cross-material generalization. More recently, DeepH-2 (Wang et al., 2024b) broadened coverage to systems involving elements primarily from the first four rows of the periodic table; however, they neglected SOC and reduced the orbital basis by omitting $f$-orbitals. While such choices help reduce computational and modeling complexity, they may limit broad applicability to the full diversity of real-world materials. Zhong et al. (2025) developed a Hamiltonian prediction model aimed at a broader range of element types, while also highlighting the challenge of achieving consistently high accuracy across diverse crystal systems. Moreover, open-source datasets with a broad and diverse collection of materials dedicated to training and validating universal Hamiltonian models across the periodic table remain scarce. Although the QH9 (Yu et al., 2023a) dataset is a well-known open-source Hamiltonian dataset, it consists of molecular systems rather than periodic material systems, and includes only structures composed of C, H, O, N, and F elements. To solve these challenges, this work presents an advanced unified deep learning framework together with a large benchmark dataset for Hamiltonian prediction, targeting broader generalization across richer classes of materials.
@@ 3 Method
As shown in Fig. 2, to effectively handle wider chemical and structural variability of materials, we develop a unified Hamiltonian prediction framework along three aspects:
@@ A.1 Electronic Structure Calculations: from Density Functional Theory to Deep Learning Methods
To address this challenge, recent approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023b; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) have adopted the deep graph learning paradigm to predict the self-consistent Hamiltonians. These methods bypass the iterative and computationally intensive matrix diagonalization steps in traditional DFT algorithms by directly predicting the final converged Hamiltonian matrix $H^{(T)}_{\alpha\beta}$ in a single forward pass. As shown in Eq. (11), the Hamiltonian matrix is inherently sparse: only pairs of atoms within a cutoff radius contribute non-zero elements. Therefore, the total number of Hamiltonian matrix elements that need to be computed scales with the number of local atomic pairs in the system, leading to a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ denotes the average number of neighboring atoms within the cutoff radius per atom. Since the atomic orbital basis functions have finite spatial support, matrix elements vanish beyond a certain inter-atomic distance. In small systems where all atoms lie within each other’s cutoff radius, $\overline{E}\sim N$, and the total number of non-zero elements scales as $\mathcal{O}(N^{2})$. However, in sufficiently large systems, $\overline{E}$ saturates to a constant determined by local geometry, making the number of non-zero Hamiltonian elements scale linearly as $\mathcal{O}(N)$. Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known. Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$ for large systems.
Recently, deep-learning approaches (Li et al., 2022; Gong et al., 2023; Yu et al., 2023b; Yin et al., 2024; Li et al., 2025; Luo et al., 2025; Yin et al., 2025) exploit the sparsity of the Hamiltonian, yielding a computational cost that scales approximately linearly with the number of non-zero matrix elements and enabling efficient, scalable prediction of quantum properties in large atomic systems. As a result, they offer a significant efficiency advantage over traditional DFT methods with computational complexity of $\mathcal{O}(TN^{3})$. This efficiency makes them particularly promising for predicting electronic structures of complex atomic systems under limited computational resources, potentially accelerating down-stream application areas like materials simulation and design.